# Efficient generation of projective modules: a motivic view

### v1

@@ 1. Introduction
Assume $R$ is a commutative ring, and $M$ is a finitely generated $R$-module. Write $\nu(M)$ for the minimal number of generators of $M$. In 1961, R. Swan posed the following problem [18, p. 272 Problem], which he attributed to J.-P. Serre.
@@
There are many variants of this kind of problem: one could drop the hypothesis that $M$ is projective, retaining only the condition that it is a finitely generated module; one could drop the hypothesis that $R$ is Noetherian, retaining only enough structure to define dimension, etc. When $R$ is Noetherian ring of Krull dimension $d$, if the localizations $M_{{\mathfrak{m}}}$ can be generated by $r$ elements for all maximal ideals $\mathfrak{m}$} of $R$, then O. Forster showed that $\nu(M)\leq r+d$ [12, Satz 1]. Shortly thereafter, R. Swan established the same result eliminating the Noetherian hypotheses on $R$: if the maximal ideal spectrum of $R$ is a Noetherian topological space of dimension $d$, and if for each maximal ideal $\mathfrak{m}$} of $R$ the module $M_{{\mathfrak{m}}}$ is generated by $r$ elements, then $\nu(M)\leq r+d$ [19, Theorem 1].
@@
One celebrated efficient generation result was established by M.P. Murthy, who observed that if $R$ is furthermore a regular $k$-algebra with $k$ an algebraically closed field, then there are geometric obstructions to efficient generation of $M$. In the context we consider, Murthy’s results imply: if $M$ is a projective $R$-module of rank $r$, then there is an associated Segre class $s_{0}(M)\in\mathrm{CH}_{0}(\operatorname{Spec}R)$ whose vanishing is necessary and sufficient for $M$ to be efficiently generated [15, Theorem 5.2 and Corollary 5.3].
@@
The key step to prove Theorem B is our identification of the first possibly nontrivial obstruction to $n$-generation of a rank $r$ module $M$ over $X=\mathrm{Spec}\,R$ a smooth affine variety, without any reference to dimension. When the ${\mathbb{A}}^{1}$-cohomological dimension of $X$ is nr+1n-r+1, there is a secondary obstruction to consider. We can identify this obstruction under suitable hypotheses.
Theorem D (cf. Propositions 21 and 37).
@@
We also identify some obstructions to expressing a symplectic module $M$ as a summand of a lower rank hyperbolic $R$-module in Theorems 44 and 46. These results depend on the parity of the (${\mathbb{A}}^{1}$-cohomological) dimension of $X$.
@@ 1.2. Preliminaries and conventions
Fix a field $k$. We write $\operatorname{Sm}_{k}$ for the category of separated, smooth $k$-schemes. Write $\operatorname{Sh}_{\operatorname{Nis}}(\operatorname{Sm}_{k})$ for the infinity category of Nisnevich sheaves of spaces on $\operatorname{Sm}_{k}$. The infinity category of motivic spaces $\operatorname{Spc}(k)$ is the full subcategory of $\operatorname{Sh}_{\operatorname{Nis}}(\operatorname{Sm}_{k})$ spanned by $\mathbb{A}^{1}$-local sheaves, i.e., those Nisnevich sheaves of spaces $\mathcal{X}$ such that the for all $U\in\operatorname{Sm}_{k}$, the map $\mathcal{X}(U)\to\mathcal{X}(U\times\mathbb{A}^{1})$ induced by projection $U\times\mathbb{A}^{1}\to U$ is an equivalence. Throughout this paper, we work in the homotopy category $\mathcal{H}(k)$ of motivic spaces over a perfect field $k$, or the pointed version. We refer the reader to [14].
@@
    Given a pointed motivic space $(\mathcal{X},x_{0})$, we write $\pi_{i}(\mathcal{X},x_{0})$ for the Nisnevich sheafification of the presheaf of groups (or sets if $i=0$) on $\operatorname{Sm}_{k}$ given by $U\mapsto\pi_{i}\mathcal{(}X(U),x_{0})$. If $\mathcal{X}\in\operatorname{Spc}_{k}$ is $\mathbb{A}^{1}$-simply connected, meaning $\pi_{0}(\mathcal{X})=*$, we write $\pi_{i}(\mathcal{X})$ for its homotopy sheaf with respect to any basepoint (cf. [14, §3.2] for discussion of homotopy sheaves).
@@
    We will frequently be concerned with strictly $\mathbb{A}^{1}$-invariant sheaves in the sense of [13, §6] or [10]. Such sheaves include the homotopy sheaves $\pi_{i}(\mathcal{X})$ for $\mathcal{X}$ a motivic space and $i\geq 2$.
@@
    We assume familiarity with affine representability results for vector bundles and symplectic vector bundles (see [13, §8.1], [8, Theorem 1], and [9, Theorem 1] for details). In particular, for $X$ a smooth affine variety over $k$, let $\operatorname{Vect}_{r}(X)$ denote the set of isomorphism classes of algebraic vector bundles of rank $r$ over the variety $X$. There is a natural bijection
@@
    where we refer the reader to [9, Page 1010] for the definition the classifying space functor $\operatorname{B}(-)$. Similarly, if $\operatorname{Vect}_{2r}^{\operatorname{Sp}}(X)$ denotes the set of isomorphism classes of symplectic rank $2r$ vector bundles over $X$, there is a natural bijection
@@
    We make extensive use of Moore–Postnikov theory in $\mathbb{A}^{1}$-homotopy theory. For the details of this obstruction theory, we refer the reader to [13, Appendix B] and [5, §6.1].
@@ 2.1. Fiber sequences involving Stiefel varieties
(2) $\mathrm{St}_{n-j}({n})\to\mathrm{Gr}_{n-j}({n})\to\mathrm{BGL}_{n},$
@@
Equation 1 follows by combining [9, Theorem 2.2.5] with [9, Lemma 2.4.1]. For Equation 2, apply [1, Proposition 3.1] to the commutative square:
@@
where the left-hand vertical map classifies direct sum of vector bundles. Taking the fiber of the induced map of horizontal fibers yields the fiber sequence Equation 1. Taking vertical fibers and applying [9, Theorem 2.2.5 and Lemma 2.4.1] to the fiber of the middle term then yields the stated result. Equation 3 follows from applying simplicial loops to the base in the sequence
@@
which is a fiber sequence by [9, Theorem 2.2.5 and Lemma 2.4.1] again. For the last sequence, the inclusions $\mathrm{GL}_{j}\subset\mathrm{GL}_{j+1}\subset\mathrm{GL}_{n}$ yields a commutative square of maps of classifying spaces
$\mathrm{BGL}_{j}$$\mathrm{BGL}_{j+1}$$\mathrm{BGL}_{n}$$\mathrm{BGL}_{n}$$\scriptstyle{\operatorname{id}}$
Then, take vertical and horizontal fibers and apply [1, Proposition 3.1]. ∎
@@
(5) $\mathrm{St}_{j}({n})$$\mathrm{Gr}_{j}({\mathbb{A}^{n}})$$\mathrm{BGL}_{j}$$(\mathbb{A}^{n-j+1}\smallsetminus 0)\times\Omega(\mathbb{A}^{n+1}\smallsetminus 0)$$\mathrm{Gr}_{j}({n})$$\mathrm{Gr}_{j}({n+1})$$\mathbb{A}^{n-j+1}\smallsetminus 0$$\mathrm{BGL}_{n-j}$$\mathrm{BGL}_{n-j+1}.$$\scriptstyle{\operatorname{pr}_{1}}$
@@
The bottom row is a fiber sequence since $\mathrm{GL}_{n-j+1}/\mathrm{GL}_{n-j}\cong\mathbb{A}^{n-j+1}\smallsetminus 0.$ For the middle second row, consider the iterated fiber of the commutative diagram:
(6) $\mathbb{A}^{n-j+1}\smallsetminus 0$$\mathbb{A}^{n+1}\smallsetminus 0$$\mathrm{Gr}_{n-j}({j})$$\mathrm{BGL}_{j}\times\mathrm{BGL}_{n-j}$$\mathrm{BGL}_{n}$$\mathrm{Gr}_{n+1-j}({n+1})$$\mathrm{BGL}_{j}\times\mathrm{BGL}_{n-j+1}$$\mathrm{BGL}_{n+1}.$}
By [13, Corollary 5.43], the top horizontal map $\mathbb{A}^{n-j+1}\smallsetminus 0\to\mathbb{A}^{n+1}\smallsetminus 0$ is nullhomotopic, so the fiber is $\mathbb{A}^{n-j+1}\smallsetminus 0\times\Omega\mathbb{A}^{n+1}\smallsetminus 0.$ We conclude by [1, Proposition 3.1]. ∎
@@
(8) $\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})\to\mathrm{Gr}^{\mathrm{Sp}}_{2n-2j}({\mathbb{A}^{2n}})\to\operatorname{B}\!\mathrm{Sp}_{2j},$
@@
(11) $\mathrm{St}^{\mathrm{Sp}}_{2j}({2n})$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({\mathbb{A}^{2n}})$$\operatorname{B}\!\mathrm{Sp}_{2j}$$(\mathbb{A}^{2(n-j+1)}\smallsetminus 0)\times(\Omega\mathbb{A}^{2(n+1)}\smallsetminus 0)$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({\mathbb{A}^{2n}})$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({\mathbb{A}^{2n+2}})$$\mathbb{A}^{2(n-j+1)}\smallsetminus 0$$\operatorname{B}\!\mathrm{Sp}_{2n-2j}$$\operatorname{B}\!\mathrm{Sp}_{2n-2j+2}$$\scriptstyle{p_{1}}$
@@ 2.2. Stabilization via weak-cellular/connectivity estimates
For each $n>j\geq 1$, the symplectic Stiefel variety $\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})$ is (2j1)(2j-1)-connected.
@@ 2.3. Computing homotopy sheaves of Stiefel varieties
$\pi_{j+1}\mathrm{St}_{n-j}({n})$ for $n>j+1$. Note that $\pi_{j}\mathrm{St}_{1}({j+1})\cong\mathrm{K}^{MW}_{j+1}$ by [13, 6.40]. The homotopy sheaf
@@
$\pi_{j+1}(\mathbb{A}^{j+1}\smallsetminus 0)\xrightarrow{\pi_{j+1}a_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j}({j})\xrightarrow{\pi_{j+1}b_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j-1}({n})\to 0$
@@
$\pi_{j+1}(\mathbb{A}^{j+1}\smallsetminus 0)\xrightarrow{\pi_{j+1}a_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j}({j})\xrightarrow{}2\mathrm{K}^{M}_{j+2}\to 0,$
@@
$\operatorname{Ker}(d_{n-j}^{j})=\operatorname{Im}(\pi_{j+1}b_{n-j}^{j})\cong 2\mathrm{K}^{M}_{j}+1.$
@@
$\pi_{2j+1}\mathrm{St}^{\mathrm{Sp}}_{2}({2j+2})\cong\pi_{2j+1}(\mathbb{A}^{2j+2}\smallsetminus 0)\cong\mathrm{K}^{MW}_{2j+1},$
@@
For any $n>j\geq 1$, $\pi_{i}\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})=0$ for $i<2j+1$ and $\pi_{2j+1}\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})\cong\mathrm{K}^{MW}_{2j+1}.$

### v2

@@ 1. Introduction
Assume $R$ is a commutative ring, and $M$ is a finitely generated $R$-module. Write $\nu(M)$ for the minimal number of generators of $M$. In 1961, R. Swan posed the following problem [28, p. 272 Problem], which he attributed to J.-P. Serre.
@@
There are many variants of this kind of problem: one could drop the hypothesis that $M$ is projective, retaining only the condition that it is a finitely generated module; one could drop the hypothesis that $R$ is Noetherian, retaining only enough structure to define dimension, etc. When $R$ is Noetherian ring of Krull dimension $d$, if the localizations $M_{{\mathfrak{m}}}$ can be generated by $r$ elements for all maximal ideals $\mathfrak{m}$} of $R$, then O. Forster showed that $\nu(M)\leq r+d$ [18, Satz 1]. Shortly thereafter, R. Swan established the same result eliminating the Noetherian hypotheses on $R$: if the maximal ideal spectrum of $R$ is a Noetherian topological space of dimension $d$, and if for each maximal ideal $\mathfrak{m}$} of $R$ the module $M_{{\mathfrak{m}}}$ is generated by $r$ elements, then $\nu(M)\leq r+d$ [29, Theorem 1].
@@
One celebrated efficient generation result was established by M.P. Murthy, who observed that if $R$ is furthermore a regular $k$-algebra with $k$ an algebraically closed field, then there are geometric obstructions to efficient generation of $M$. In the context we consider, Murthy’s results imply: if $M$ is a projective $R$-module of rank $r$, then there is an associated Segre class $s_{0}(M)\in\mathrm{CH}_{0}(\operatorname{Spec}R)$ whose vanishing is necessary and sufficient for $M$ to be efficiently generated [24, Theorem 5.2 and Corollary 5.3].
@@
The key step to prove Theorem B is our identification of the first possibly nontrivial obstruction to $n$-generation of a rank $r$ module $M$ over $X=\mathrm{Spec}\,R$ a smooth affine variety, without any reference to dimension. When the ${\mathbb{A}}^{1}$-cohomological dimension of $X$ is greater than nr+1n-r+1, there is a secondary obstruction to consider. We can identify this obstruction under suitable hypotheses.
Theorem D (cf. Propositions 21 and 39).
@@
We also identify some obstructions to expressing a symplectic module $M$ as a summand of a lower rank hyperbolic $R$-module in Theorems 46 and 48. These results depend on the parity of the (${\mathbb{A}}^{1}$-cohomological) dimension of $X$.
1.1. Related literature
As already indicated, there are many variants of 1 above. While this paper presents a motivic approach to efficient generation of projective modules and symplectic modules, work by many authors studies efficient generation in other cases. For example, Eisenbud and Evans proved a strengthening of the Forster–Swan theorem that considers how different generating sets are related [15, Theorem B]. Extensive results have been obtained for ideals in polynomial rings in [26, 21] and [13], and for ideals in overrings of polynomial rings in [14]. More recently, efficient generation results were proven for ideals in geometric subrings of polynomial rings in [12] and for ideals in discrete Hodge algebras in [20] and [19].
@@ 1.3. Preliminaries and conventions
Fix a field $k$. We write $\operatorname{Sm}_{k}$ for the category of separated, smooth $k$-schemes. Write $\operatorname{Sh}_{\operatorname{Nis}}(\operatorname{Sm}_{k})$ for the infinity category of Nisnevich sheaves of spaces on $\operatorname{Sm}_{k}$. The infinity category of motivic spaces $\operatorname{Spc}(k)$ is the full subcategory of $\operatorname{Sh}_{\operatorname{Nis}}(\operatorname{Sm}_{k})$ spanned by $\mathbb{A}^{1}$-local sheaves, i.e., those Nisnevich sheaves of spaces $\mathcal{X}$ such that for all $U\in\operatorname{Sm}_{k}$, the map $\mathcal{X}(U)\to\mathcal{X}(U\times\mathbb{A}^{1})$ induced by projection $U\times\mathbb{A}^{1}\to U$ is an equivalence. Throughout this paper, we work in the homotopy category $\mathcal{H}(k)$ of motivic spaces over a perfect field $k$, or the pointed version. We refer the reader to [23].
@@
    Given a pointed motivic space $(\mathcal{X},x_{0})$, we write $\pi_{i}(\mathcal{X},x_{0})$ for the Nisnevich sheafification of the presheaf of groups (or sets if $i=0$) on $\operatorname{Sm}_{k}$ given by $U\mapsto\pi_{i}\mathcal{(}X(U),x_{0})$. If $\mathcal{X}\in\operatorname{Spc}_{k}$ is $\mathbb{A}^{1}$-simply connected, meaning $\pi_{0}(\mathcal{X})=*$ and $\pi_{1}(\mathcal{X})=*$, we write $\pi_{i}(\mathcal{X})$ for its homotopy sheaf with respect to any basepoint (cf. [23, §3.2] for discussion of homotopy sheaves).
@@
    We will frequently be concerned with strictly $\mathbb{A}^{1}$-invariant sheaves in the sense of [22, §6] or [11]. Such sheaves include the homotopy sheaves $\pi_{i}(\mathcal{X})$ for $\mathcal{X}$ a motivic space and $i\geq 2$.
@@
    We assume familiarity with affine representability results for vector bundles and symplectic vector bundles (see [22, §8.1], [9, Theorem 1], and [10, Theorem 1] for details). In particular, for $X$ a smooth affine variety over $k$, let $\operatorname{Vect}_{r}(X)$ denote the set of isomorphism classes of algebraic vector bundles of rank $r$ over the variety $X$. There is a natural bijection
@@
    where we refer the reader to [10, Page 1010] for the definition the classifying space functor $\operatorname{B}(-)$. Similarly, if $\operatorname{Vect}_{2r}^{\operatorname{Sp}}(X)$ denotes the set of isomorphism classes of symplectic rank $2r$ vector bundles over $X$, there is a natural bijection
@@
    We make extensive use of Moore–Postnikov theory in $\mathbb{A}^{1}$-homotopy theory. For the details of this obstruction theory, we refer the reader to [22, Appendix B] and [5, §6.1].
@@ 2.1. Fiber sequences involving Stiefel varieties
(2) $\mathrm{St}_{n-j}({n})\to\mathrm{Gr}_{j}({n})\to\mathrm{BGL}_{j},$
@@
Equation 1 follows by combining [10, Theorem 2.2.5] with [10, Lemma 2.4.1], which we also appeal to for the other items in combination with additional arguments. For Equation 2, apply [1, Proposition 3.1] to the commutative square:
@@
where the left-hand vertical map classifies direct sum of vector bundles: we take the fiber of the induced map of horizontal fibers. Equation 3 follows from applying simplicial loops to the base in the sequence
@@
which is a fiber sequence by [10, Theorem 2.2.5 and Lemma 2.4.1] again. For the last sequence, the argument is analogous Equation 2: the inclusions $\mathrm{GL}_{j}\subset\mathrm{GL}_{j+1}\subset\mathrm{GL}_{n}$ yields a commutative square of maps of classifying spaces
$\mathrm{BGL}_{j}$$\mathrm{BGL}_{j+1}$$\mathrm{BGL}_{n}$$\mathrm{BGL}_{n},$$\scriptstyle{\operatorname{id}}$
and we apply [1, Proposition 3.1]. ∎
@@
(5) $\mathrm{St}_{j}({n})$$\mathrm{Gr}_{j}({n})$$\mathrm{BGL}_{j}$$(\mathbb{A}^{n-j+1}\smallsetminus 0)\times\Omega(\mathbb{A}^{n+1}\smallsetminus 0)$$\mathrm{Gr}_{j}({n})$$\mathrm{Gr}_{j}({n+1})$$\mathbb{A}^{n-j+1}\smallsetminus 0$$\mathrm{BGL}_{n-j}$$\mathrm{BGL}_{n-j+1}.$$\scriptstyle{\operatorname{pr}_{1}}$
@@
The bottom row is a fiber sequence since $\mathrm{GL}_{n-j+1}/\mathrm{GL}_{n-j}\cong\mathbb{A}^{n-j+1}\smallsetminus 0.$ For the second row, consider the iterated fiber of the commutative diagram:
(6) $\mathbb{A}^{n-j+1}\smallsetminus 0$$\mathbb{A}^{n+1}\smallsetminus 0$$\mathrm{Gr}_{j}({n})$$\mathrm{BGL}_{j}\times\mathrm{BGL}_{n-j}$$\mathrm{BGL}_{n}$$\mathrm{Gr}_{j}({n+1})$$\mathrm{BGL}_{j}\times\mathrm{BGL}_{n-j+1}$$\mathrm{BGL}_{n+1}.$}
By [22, Corollary 5.43], the top horizontal map $\mathbb{A}^{n-j+1}\smallsetminus 0\to\mathbb{A}^{n+1}\smallsetminus 0$ is nullhomotopic, so the fiber is $\mathbb{A}^{n-j+1}\smallsetminus 0\times\Omega\mathbb{A}^{n+1}\smallsetminus 0.$ We conclude by [1, Proposition 3.1]. ∎
@@
(8) $\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})\to\mathrm{Gr}^{\mathrm{Sp}}_{2j}({2n})\to\operatorname{B}\!\mathrm{Sp}_{2j},$
@@
(11) $\mathrm{St}^{\mathrm{Sp}}_{2j}({2n})$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({2n})$$\operatorname{B}\!\mathrm{Sp}_{2j}$$(\mathbb{A}^{2(n-j+1)}\smallsetminus 0)\times(\Omega\mathbb{A}^{2(n+1)}\smallsetminus 0)$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({2n})$$\mathrm{Gr}^{\mathrm{Sp}}_{2j}({2n+2})$$\mathbb{A}^{2(n-j+1)}\smallsetminus 0$$\operatorname{B}\!\mathrm{Sp}_{2n-2j}$$\operatorname{B}\!\mathrm{Sp}_{2n-2j+2}$$\scriptstyle{p_{1}}$
@@ 2.2. Stabilization via weak-cellular/connectivity estimates
For each $n>j\geq 1$, the symplectic Stiefel variety $\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})$ is $2j$-connected.
@@ 2.3. Computing homotopy sheaves of Stiefel varieties
$\pi_{j+1}\mathrm{St}_{n-j}({n})$ for $n>j+1$. Note that $\pi_{j}\mathrm{St}_{1}({j+1})\cong\mathrm{K}^{MW}_{j+1}$ by [22, 6.40]. The homotopy sheaf
@@
$\pi_{j+1}(\mathbb{A}^{j+1}\smallsetminus 0)\xrightarrow{\pi_{j+1}a_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j}({n})\xrightarrow{\pi_{j+1}b_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j-1}({n})\to 0$
@@
$\pi_{j+1}(\mathbb{A}^{j+1}\smallsetminus 0)\xrightarrow{\pi_{j+1}a_{n-j}^{j}}\pi_{j+1}\mathrm{St}_{n-j}({n})\xrightarrow{}2\mathrm{K}^{M}_{j+2}\to 0,$
@@
$\operatorname{Ker}(d_{n-j}^{j})=\operatorname{Im}(\pi_{j+1}b_{n-j}^{j})\cong 2\mathrm{K}^{M}_{j+2}.$
@@
$\pi_{2j+1}\mathrm{St}^{\mathrm{Sp}}_{2}({2j+2})\cong\pi_{2j+1}(\mathbb{A}^{2j+2}\smallsetminus 0)\cong\mathrm{K}^{MW}_{2j+2},$
@@
For any $n>j\geq 1$, $\pi_{i}\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})=0$ for $i<2j+1$ and $\pi_{2j+1}\mathrm{St}^{\mathrm{Sp}}_{2n-2j}({2n})\cong\mathrm{K}^{MW}_{2j+2}.$